[Hallc_running] [New Logentry] Hal C RC Update: Sunday 8/26

gaskelld at jlab.org gaskelld at jlab.org
Sun Aug 26 08:40:08 EDT 2018

Logentry Text:
As yesterday, the goal for today is still to deliver CW beam at 4 pass to Hall C. Problems with an FSD related to the beam dump and an ion pump in the South Linac slowed progress yesterday.   Overnight and this morning, we have seen periods of tune beam in the hall, but optics work is still ongoing.

The plan for today is the same as yesterday. Once we have CW, the SHMS steering test plan will be executed, and then we will move on to the KAON-LT run plan.

There will be no RC meeting today.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3591433
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180826/93931fa3/attachment.html>

More information about the Hallc_running mailing list